# LoDIP: Low light phase retrieval with deep image prior

### v1

@@ V Conclusion and Future Work
Emerging deep learning methods present several opportunities to improve upon and expand the scope of computational imaging techniques that require processing steps framed as mathematical optimization problems. The LoDIP approach, introduced in this work, integrates the optimization framework of untrained neural network priors with a precise forward model inspired by the in-situ CDI setup. This combination aims to recover sample images from measured far-field diffraction patterns, particularly in situations where well-established iterative algorithms struggle to achieve proper convergence. We have demonstrated that LoDIP yields an improvement in reconstruction across multiple metrics and samples, especially under low-dose conditions. Moreover, it does not require labeled training data, complex tuning of hyper-parameters, or supervision from a specialist in phase retrieval algorithms. We tested LoDIP for multiple samples and static regions including synthetic data generated from google images, realistic biological cells and experimental data with known and unknown static regions. Our experiments demonstrate LoDIP’s versatility, showcasing its successful application to a wide variety of samples. Furthermore, we showed that LoDIP outperforms state-of-the-art methods in obtaining a denoised reconstructed image, while preserving resolution. We anticipate applications of the LoDIP method to X-ray imaging of dose-sensitive samples of importance in a variety of fields, such as organic semiconductors relevant to modern perovskite solar cells and battery materials, as well as biological samples concerning cellular interactions and life cycles.
@@ Acknowledgments
The work is supported by STROBE: A National Science Foundation Science & Technology Center under grant number DMR 1548924. E.N. is also partially supported by the Simons Postdoctoral program at IPAM and NSF DMS 1925919 and by AFOSR MURI FA9550-21-1-0084. The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported within this paper.
The authors thank Jianwei (John) Miao (UCLA) and Stanley J. Osher (UCLA) for their valuable feedback while completing this project.

### v2

@@ V Conclusion and Future Work
Emerging deep learning methods present several opportunities to improve upon and expand the scope of computational imaging techniques framed as mathematical optimization problems. The LoDIP approach, introduced in this work, integrates the optimization framework of implicit generative priors with physical constraints from a modified CDI setup. This combination results in high-SNR reconstructions while preserving resolution at low photon counts in strong presence of noise where traditional methods are known to struggle. Experiments demonstrate that LoDIP’s performance is consistent across different natural images, realistic biological cells and experimental data. We expect the LoDIP method to find applications in X-ray imaging of dose-sensitive samples across diverse fields including organic semiconductors and biological specimens.
@@ Acknowledgments
This work was supported by STROBE, a National Science Foundation Science & Technology Center, under grant number DMR-1548924, and the U.S. Air Force Office of Scientific Research Multidisciplinary University Research Initiative (MURI) program under Award No. FA9550-23-1-0281. E.N. was partially supported by the Simons Postdoctoral program at IPAM, NSF DMS-1925919, and NSF-2331033. The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported in this paper.